OpenStudy (anonymous):

Which equation represents this line in point-slope form? The points are (-3,4) and (3,-4) A) (y+4)=4/3(x+3) B) (y−4)=−4/3(x+3) C) y=−3/4x D) 4y−3x=0

OpenStudy (muzzack):

m=-4-4/3+3 m=-8/6 m=-4/3 ------- y-4=-4/3(x+3)

OpenStudy (muzzack):

B is ur answer

OpenStudy (anonymous):

Thank you!!